how do you graph the linear function represented by the equation y=240-10x

Answer by mananth(16949)

y=240-10x
generarte values for y from different values of x
when x= 30 , y= 140 (10,-60)
when x=20,y = 40 (20,40)
when x= 25,y=-10 (25,-10)